Fermi National Laboratory
 

Operation Summary for September 21, 2011 - September 23, 2011

Wednesday September 21

The day shift began with the Tevatron (TEV) in shot setup, with Operations monitoring the Recycler stash (50.8E10), with the Antiproton Source (Pbar) stacking (68.4E10), with the Main Injector (MI), Linac, and Booster operating normally, with NuMI off for a target change, with the MiniBooNE experiment taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for the next store.

During the day shift, Operations established store 9142 at 8:33 AM with an average initial luminosity (AIL) of 350.68E30. Operations reported at 9:11 AM that they had received permission to send beam to a new Meson FTBF experiment T-1017. Operations established beam to T-1017 at 9:36 AM. MI and Instrumentation personnel began a test of new software for the MI BPMs . MI personnel began a coalescing study at 3:01 PM. The MI BPM system test ended at 3:17 PM and the system was put back to normal.

During the evening shift, Booster personnel reported at 4:33 PM that they would be conducting a study that would halt beam for 10 minutes. Operations disabled beam to T-1017 at 6:01 PM. At 7:35 PM, a Linac RF station (LRF3) tripped off for the first of four times in two hours and 45 minutes; operators called in a technician after the last trip, but then the station began operating normally. Operations terminated store 9142 at 9:53 PM, put the TEV into HEP shot setup, and established store 9143 at 11:59 PM with an AIL of 343.25E30. Operators reported short downtimes due to a Linac quadrupole magnet power supply (L:QPS413), two Linac Klystron RF stations (KRF1 and KRF6), and a Booster RF station (BRF7).

Thursday September 22

The midnight shift began with Operations monitoring store 9143 (313.86E30) and the stash (43.5E10), with Pbar stacking (98.4E10), and with MiniBooNE taking beam.

During the midnight shift, at 5:38 AM, operators reported at I- Source vacuum valve closed; the vacuum recovered and beam returned. Operations established beam to T-1017 at 6:45 AM.

During the day shift, the Neutron Therapy Facility (NTF) began tuning their beam line in preparation to treat one patient later in the morning. Operators reported that NuMI personnel had completed the installation of their new NuMI target; at 12:43 PM they began low intensity scans. Operations terminated store 9143 at 1:08 PM. Operators prepared D0 and CDF for accesses; the accesses ended and operators put the TEV through a dry squeeze at 2:54 PM. NuMI personnel completed the low intensity target scans at 3:32 PM. Operators prepared the TEV for a study using two high intensity proton bunches; operators put the TEV through a wet squeeze at 3:57 PM.

During the evening shift, TEV personnel completed their study. Operations put the TEV into shot setup at 4:22 PM and established store 9146 at 5:44 PM with an AIL of 347.3E30. Operations disabled beam to T-1017 at 6:08 PM. Operators reported at 7:42 PM a MI-8 vacuum burst; everything reset. At 8:49 PM, three of the four vacuum pumps for the I- Source tripped off; everything reset.

Friday September 23

The midnight shift began with Operations monitoring store 9146 (133.55E30) and the stash (193.0E10), with Pbar stacking (20.7E10), and with MiniBooNE taking beam.

During the midnight shift, operators reported at 12:33 AM that BRF1 tripped off for the first of three times in 16 minutes; it reset. At 4:31 AM, the MI-52 LCW pump #3 tripped off; operators tried to restart it, but it tripped off again. Operators investigated and turned the pump on locally and everything seemed normal. Operators contacted a Fluids person who instructed them to bring on pump #1 in its place. Operations terminated store 9146 at 7:56 AM.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store. The Run Coordinator will allow Booster and MI personnel to conduct studies. NuMI personnel reported finding a number of cooling water leaks associated with the new target, but not in the target. They will repair the leaks today and then watch the cooling water system through today; this will delay their taking beam until Saturday morning (9/24/11).


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 09/23/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory